# Advancing Multi-Modal Sensing through Expandable Modality Alignment

### v1

@@ 8.1. Evaluation on HAR
The rationale that these MLLMs seems supporting sensing modalities, but struggle to comprehend IMU data and manage HAR tasks, is their training only on the Ego4D dataset (Grauman et al., 2022). Without sufficient training, these models are restricted to trained data, limiting their cross-domain capabilities. Furthermore, these MLLMs are unable to be trained on other sensing datasets due to data scarcity and absence of techniques like pre-trained modality tower and the expandable architecture, which are introduced in Babel.
8.1.5. Ablation Study and Growth Orders.
The techniques, including the pre-trained encoders, expandable network architecture and adaptive training strategy, are all essential for constructing Babel. Particularly, Without pre-trained modality tower, training wouldn’t converge due to limited samples. On UTD-MHAD (Chen et al., 2015), without prototype network, the previously aligned modality would drop about averagely 44.7% relatively after introducing a new modality. Without adaptive training, the overall performance would decrease by up to 7.2%.
@@
Thanks to the techniques proposed in Babel, the order of modality growth does not significantly influence the end-to-end performance once the training is sufficient. To evaluate this, we devise four network growth sequences according to different heuristics: (i) random order; (ii) alignment from the most robust to weakest modality (skeleton, video, LiDAR, IMU, Wi-Fi, mmWave); (iii) alignment based on data diversity, taking into account the number of actions, subjects, and scenes of the used datasets; (iv) alignment based on the data amount of used datasets, organized from largest to smallest. As shown in Table 5, growth order doesn’t significantly affect the performance. For instance, with different growth orders, the performance on IMU and Wi-Fi modality varies less than 3% and 2%, respectively. This highlights Babel’s robustness.
@@
8.1.6. System Overhead
The pre-trained hexa-modal Babel takes around 1.1GB on the disk, including pre-trained encoders, concept alignment modules and the prototype network. Babel takes 1.4-9.92GB memory, depending on the selected modalities, using the FP32 precision. We evaluate the Babel’s inference latency on one Nvidia A100 GPU. The exact inference latency depends on the selected modalities. For instance, per each sample, i.e., a sequence of sensing data spanning 3-4 seconds, Babel takes 98.6ms for IMU modality, 206.7ms for LiDAR, For skeleton, WiFi, and mmWave, Babel only takes around 130ms. When fusing modalities, modality towers could be executed in parallel. Babel takes 265ms for Wi-Fi and video, 138ms for IMU and skeleton together.
@@ 8.2.1. Cross-modality retrieval
9. Related Work
Multi-modal Sensing. The development of multi-modal sensing networks is an emerging research area. Recently Cosmo (Ouyang et al., 2022) pioneered the application of contrastive fusion learning in multi-modal sensing, incorporating RGB, depth and IMU modalities. MESEN (Xu et al., 2023) employs multi-modal contrastive learning to improve the performance of singular-modal sensing. In contrast to Cosmo and MESEN, Babel, to the best of our knowledge, is the first expandable pre-trained network for multi-modal sensing, which seamlessly facilitates the alignment of additional modalities from diverse, scattered datasets. The concept of multi-modal sensing is extensively employed across a wide array of applications. For instance, (Liu et al., 2021b) integrates RFID and RGB for recognizing human-object interactions. (Wang et al., 2022) leverages LiDARs, cameras, and IMU and GNSS devices worn by animals to recognize animal behavior. To locate target individuals, (Liu et al., 2021a) utilizes Wi-Fi Fine Timing Measurements and IMU data to associate individuals in a video with a matched query ID. GaitVibe+ (Dong et al., 2022) enhances structural vibration-based footstep localization using temporary cameras and vibration sensors for in-home gait analysis. (Girod and Estrin, 2001) presents an acoustic and camera sensing system that ameliorates range estimation for applications in robotics and other domains. These applications could potentially benefit through Babel.
Modality Alignment and MLLMs. Contrastive Learning (CL) is widely used for modality alignment, particularly in visual and linguistic modalities (Radford et al., 2021; Zhai et al., 2022; Rosenfeld et al., 2022). Despite the prevalence of CL, its implementation in multi-modal sensing introduces significant challenges. To overcome data scarcity, we have introduced several techniques in Babel, which are essential for aligning multiple sensing modalities. There is also a growing trend of aligning a broader range of modalities. For example, Meta-Transformer (Zhang et al., 2023a) uses a unified frozen visual encoder to derive representations across 12 different modalities. ImageBind (Girdhar et al., 2023) combines six modalities using only image-paired data through contrastive learning. However, these approaches provide limited support for sensing modalities, and their performance is suboptimal, as demonstrated in Table 4. This is due to their lack of sufficient training without an expandable architecture like Babel. Discussions have also been held regarding the ability of LLMs to reason about the real physical world (Xu et al., 2024). (Zhang et al., 2023a; Girdhar et al., 2023) usually serve as multi-modal encoders to integrate with LLMs for understanding sensing signals. For instance, OneLLM (Han et al., 2023) employs Meta-Transformer (Zhang et al., 2023a), while M4 (Yuan et al., 2024) and AnyMAL (Moon et al., 2023) use ImageBind (Girdhar et al., 2023). Babel, as a scalable pre-trained network, could potentially advance these MLLMs by better understanding more sensing modalities, which is our future work.
10. Conclusion and Future Work
We present Babel, a expandable modality alignment framework designed for sensing applications. The pre-trained Babel has been proficiently aligned with six prevalent sensing modalities, IMU, skeleton, video, Wi-Fi, LiDAR, and mmWave. Babel demonstrated the superior performance for HAR tasks across various datasets compared to an array of baselines. As Babel is a scalable network, we call for the community to further enhance and align additional helpful modalities into Babel.

### v2

@@ 8.1. Evaluation on HAR
The rationale that these MLLMs seems supporting sensing modalities, but struggle to comprehend IMU data and manage HAR tasks, is their training only on the Ego4D dataset (Grauman et al., 2022). Without sufficient training, these models are restricted to trained data, limiting their cross-domain capabilities. Furthermore, these MLLMs are unable to be trained on other sensing datasets due to data scarcity and absence of techniques like pre-trained modality tower and the expandable architecture, which are introduced in Babel. MLLMs could be improved by incooperating Babel as the sensing modality encoders, which would be discussed in §8.2
8.1.5. Growth Orders
@@
Thanks to the prototype network architecture and the adaptive training strategy proposed in Babel, the order of modality growth does not significantly influence the end-to-end performance once the training is sufficient. To evaluate this, we devise four network growth sequences according to different heuristics: (i) random order: the modalities are aligned in the sequence of IMU, skeleton, video, Wi-Fi, mmWave, and LiDAR. (ii) alignment from the most robust to weakest modality (skeleton, video, LiDAR, IMU, Wi-Fi, mmWave); (iii) alignment based on data diversity, taking into account the number of actions, subjects, and scenes of the used datasets; The sequence follows skeleton, mmWave, LiDAR, IMU, and Wi-Fi. (iv) alignment based on the data amount of used datasets, organized from largest to smallest, proceeding from skeleton, mmWave, Wi-Fi, LiDAR, and IMU. Using these different growth orders, we train Babel and evaluate the end-to-end classification accuracy on the downstream tasks. As shown in Table 5, growth order doesn’t significantly affect the performance. For instance, with different growth orders, the performance on IMU and Wi-Fi modality varies less than 3% and 2%, respectively. This highlights Babel’s robustness.
8.1.6. Ablation
The techniques proposed in Babel, including the pre-trained encoders, expandable network architecture and adaptive training strategy, are all essential for constructing Babel. Particularly, Without pre-trained modality tower, training wouldn’t converge due to limited samples. On UTD-MHAD (Chen et al., 2015), without prototype network, the previously aligned modality would drop about averagely 44.7% relatively after introducing a new modality. Without adaptive training, the overall performance would decrease by up to 7.2%.
@@ 8.1.6. Ablation
Figure 8. Per-sample inference latency breakdown for each modality in Babel evaluated on NVidia A100 GPU.
8.1.7. System Overhead.
In Babel, we demonstrates efficiency by introducing minimal additional system overhead. Fig. 8 illustrates the breakdown of inference latency for each modality in Babel, evaluated on a NVidia A100 GPU for an activity sample. As depicted, the overhead introduced by Babel, specifically the alignment module, is very limited compared to the pre-trained encoders. For example, in the IMU modality, the alignment module requires approximately 10.1 ms, while the pre-trained encoder, LIMU-BERT, takes 182.4 ms to process a fixed window of IMU signals. On average, the alignment module incurs only about an 8% increase in inference overhead based on our evaluation. The overhead of the alignment module varies across modalities due to the use of different MLP layer configurations to accord with each modality’s encoder. Aside from the alignment module, the prototype network is shared among modalities, with its overhead being negligible, less than 1 ms. When applying multiple modalities to downstream tasks, each modality tower can be parallelized to hide multi-modal sensing latency.
The pre-trained weights of Babel occupy approximately 1.1GB on disk, encompassing pre-trained encoders for all modalities, concept alignment modules, and the prototype network. Depending on the selected modalities, Babel requires 1.4-9.92GB of memory using FP32 precision.
@@ 8.2.1. Cross-modality retrieval
9. Discussion and Future Work
While Babel demonstrates promising results in aligning multiple sensing modalities for HAR tasks, several directions remain for our future exploration. Firstly, the framework could be extended to other sensing applications including localization, gesture detection (Sun et al., 2020), and autonomous navigation. Secondly, Our preliminary exploration of integrating Babel with LLMs demonstrates the potential for enhanced sensing comprehension. The current approach of aligning sensing embeddings with video encoders in MLLMs. Future research should investigate direct integration methods that can map sensing features into LLM’s native representation spaces without requiring intermediate alignment steps. Our case studies in sensing imaging and LLM integration could be further benefit from more state-of-the-art diffusion models (Wu et al., 2023; Xia et al., 2024; Hu et al., 2024) and LLMs (Lin et al., 2023; Zhang et al., 2024; Li et al., 2024).
10. Conclusion
We present Babel, a expandable modality alignment model designed for sensing applications. The pre-trained Babel has been proficiently aligned with six prevalent sensing modalities, IMU, skeleton, video, Wi-Fi, LiDAR, and mmWave. Babel demonstrated the superior performance for HAR tasks across various datasets compared to an array of baselines. As Babel is a scalable network, we call for the community to further enhance and align additional helpful modalities into Babel.